( \begin{matrix}a_{11} & a_{12} \\ a_{21} & a_{22} \end{matrix} )

f(x) = \{ \begin{matrix}1 & if x > 0. \\ 0 & if x = 0. \\ −1 & if x < 0. \end{matrix}

f(x) = \{ \begin{matrix}1 & if x > 0. \\ 0 & if x = 0. \\ −1 & if x < 0. \end{matrix}

( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ≠ ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} ) ( \begin{matrix}1 \\ 2 \\ 3 \end{matrix} )